OpenStudy (anonymous):

Help Please!! *I'll fan and give a medal!* Thanks. I have to evaluate the equation. a=3 g^2 x a^3= ? Answers: 162 486 729 2,187

OpenStudy (anonymous):

what is value of g?

OpenStudy (anonymous):

It doesn't say. The only Value it gives is for the Value of a.

OpenStudy (anonymous):

a=3 g^2 x a^3= ? there is wrong you have to put how much it equal or give value for g?

OpenStudy (anonymous):

g^2 x 27= ? it can slove by g= \[\sqrt{\frac{ ? }{ 27 }} \]

OpenStudy (anonymous):

mmk. I gotta find out what it all equals though but without knowing the Value of G there is no way to get a right answer because you could plug in number after number and it could give you every number that's an option. Thanks for the help though. :)